### v1

# Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs

Abstract
This work investigates the ability of open Large Language Models (LLMs) to predict citation intent through in-context learning and fine-tuning. Unlike traditional approaches that rely on pre-trained models like SciBERT, which require extensive domain-specific pretraining and specialized architectures, we demonstrate that general-purpose LLMs can be adapted to this task with minimal task-specific data. We evaluate twelve model variations across five prominent open LLM families using zero, one, few, and many-shot prompting to assess performance across scenarios. Our experimental study identifies the top-performing model through extensive experimentation of in-context learning-related parameters, which we fine-tune to further enhance task performance. The results highlight the strengths and limitations of LLMs in recognizing citation intents, providing valuable insights for model selection and prompt engineering. Additionally, we make our end-to-end evaluation framework and models openly available for future use.
Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs
Paris Koloveas IMSI, ATHENA RC / Athens, GR University of the Peloponnese / Tripolis, GR pkoloveas@athenarc.gr                        Serafeim Chatzopoulos IMSI, ATHENA RC / Athens, GR schatz@athenarc.gr
Thanasis Vergoulis IMSI, ATHENA RC / Athens, GR vergoulis@athenarc.gr                        Christos Tryfonopoulos University of the Peloponnese / Tripolis, GR trifon@uop.gr
1 Introduction
Citations are references of research articles to external sources of information included to support claims, provide context, criticize, or acknowledge prior work. Although their primary function is to inform and redirect the reader, citations are also frequently used for other purposes, such as serving as proxies of scientific impact in various types of analysis (Kanellos et al., 2021). In such use cases, understanding the exact intent of a citation is crucial, as not all types of citations should be considered. For instance, while measuring the impact of an article, citations that criticize the work should generally not be considered as contributing to its impact.
Predicting citation intent based on its context (i.e., the sentences in the manuscript that accompany the citation) and other related information, has become an important classification problem to support the aforementioned use cases. Existing approaches have traditionally relied on linguistic features (Jurgens et al., 2018), machine learning methods (Teufel et al., 2006), and, in recent years, on pre-trained language models (PLMs) such as SciBERT (Beltagy et al., 2019) that require large scientific datasets such as (Lo et al., 2020) and task-specific architectures. In contrast, this work is the first to explore the potential of open, general-purpose Large Language Models (LLMs) to accurately identify citation intent without using any scientific-trained PLMs, evaluating their effectiveness through in-context learning and fine-tuning on minimal task-specific data.
Large Language Models (LLMs) are advanced natural language processing systems trained on extensive text corpora to perform a wide range of language tasks. Unlike traditional pre-trained language models (PLMs), LLMs are general-purpose models capable of adapting to new tasks with minimal additional training. Their ability to process and generate coherent text across diverse contexts makes them particularly suitable for tasks like citation intent classification, where understanding nuanced language patterns is essential.
In-context learning is a paradigm that allows language models to learn tasks given only a few examples in the form of demonstration (Brown et al., 2020; Dong et al., 2024). In-context learning is particularly suited to citation intent classification, as it allows the models to leverage their contextual understanding of language to make accurate predictions.
In our study, we conduct an extensive experimental analysis of $12$ general-purpose instruction-tuned LLMs from $5$ model families, on two widely used datasets for this task. In this context, we analyze the impact of multiple parameters on model performance and identify the optimal configurations and best-performing models for our problem. These experiments address the following research questions:
  • RQ1: How well can pre-trained LLMs perform on citation intent classification without task-specific training?
  • RQ2: What are the differences in performance between open LLMs of varying parameter counts?
  • RQ3: How do different prompting-related parameters affect model performance?
Furthermore, we take the analysis a step further by selecting the top-performing cases from the aforementioned experiments for Supervised Fine-Tuning (SFT) on our datasets, allowing us to evaluate their optimal performance. This essentially addresses an additional research question:
  • RQ4: How much does supervised fine-tuning with task-specific training affect the performance of the instruction-tuned models?
Our experiments aim to contribute to the broader research community by guiding prompt engineering and model selection strategies for similar tasks. To further assist researchers, we openly [^1] https://github.com/athenarc/CitationIntentOpenLLM provide our complete testing suite and evaluation results, allowing seamless integration of Hugging Face models and easy configuration adjustments. We also publish the weights and checkpoints of the fine-tuned models [^2] https://huggingface.co/collections/sknow-lab/citationintentllm-67b72f1d5ca6113f960dba04 presented in Section 4.
2 Methodology
In this section, we elaborate on the evaluation methodology followed in our experimental study. We first provide an overview of the used models (Section 2.1), datasets (Section 2.2), and configuration parameters (Section 2.3). Finally, in Section 2.4, we present the technical specifications of the system we used for the experimentation.
2.1 Models
For our experiments, we selected the Instruction-tuned versions of five prominent open-weight model families: LLaMA (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), and Qwen (Bai et al., 2023; Qwen et al., 2025). A key challenge was performing the experiments on commodity hardware, both due to computational limitations and to demonstrate that citation intent classification can be executed efficiently with limited resources. This influenced the selection of models, which range in size from small ($1$B parameters) to medium-sized ($32$B parameters).
Since in-context learning inherently increases the number of tokens to each prompt (particularly in the many-shot scenario), we opted for a lower cutoff of $8,192$ tokens in the context length. This ensures that all selected models could process the longest prompts in our experiments without truncation.
To reduce the memory footprint and computational requirements of our evaluation, we utilize the 8-bit (Q8) quantized versions of the models. This approach significantly reduces memory usage without compromising performance. Quantization involves converting model parameters from 16-bit floating-point precision to 8-bit integers, enabling more efficient computation while preserving expressive power and accuracy (Jurafsky and Martin, 2025).
The model variations used for our evaluation were the following (a more detailed view is available in Appendix C):
  • Llama 3 & 3.1 (8B), Llama 3.2 (1B, 3B)
  • Mistral Nemo (12B)
  • Phi 3 Medium (14B), Phi 3.5 Mini (3.8B)
  • Gemma 2 (2B, 9B, 27B)
  • Qwen 2 (7B), Qwen 2.5 (14B)
2.2 Datasets
For our experiments, we used two datasets:
  • SciCite (Cohan et al., 2019) consists of approximately $11,000$ citation strings annotated with three classes: Background Information, Method, and Results Comparison.
  • ACL-ARC (Jurgens et al., 2018) contains 2,000 citation strings classified into six categories: Background, Motivation, Uses, Extends, Compares or Contrasts, and Future.
These two datasets are widely used in citation intent classification research. The SciCite dataset is larger and more diverse, while the ACL-ARC dataset offers a more granular classification scheme.
2.3 Configuration Parameters
In this section, we describe the configuration parameters of the examined models, including prompting methods, system prompts, query templates, example methods, and temperature settings.
2.3.1 In-context Learning Methods
To evaluate model performance, we applied four prompting methods (PM) by following the In-Context Learning paradigm (Brown et al., 2020): Zero-shot (no examples), One-shot (a single example per class), Few-shot ($5$ examples per class), and Many-shot ($10$ examples per class).
The aforementioned prompting methods were selected to evaluate whether model performance improves as the number of examples increases, and to identify any saturation point where additional examples yield diminishing or negative returns. This aligns with prior literature on in-context learning methods, where these specific configurations have been extensively studied (Dong et al., 2024). Incorporating examples directly into the prompts allows models to better understand the task and the expected output format, which is especially important for citation intent classification, where the citation context is key to determining the correct label.
2.3.2 System Prompts
System prompts (SP) play a critical role in our experimental design, as they set the context and expectations for the model’s behavior. They guide the model’s attention on citation intent classification, providing task-specific context, class definitions, and guidelines to ensure outputs are aligned with the evaluation requirements.
We explored three distinct variations of system prompts. The first (SP1) was a simple, intuitive instruction, serving as a natural starting point for introducing the task. The second prompt (SP2) adopted a structured approach inspired by the CO-STAR framework (Teo, 2023), which organizes prompts into six key components. In particular, Context provides background information to help the model understand the scenario, while Objective clearly defines the task to direct its focus. Style, Tone, and Audience shape the model’s writing style, sentiment, and intended readership, respectively. Finally, Response defines the expected format to ensure clarity and downstream processing.
Since our task requires a single citation label as output, we excluded stylistic components (i.e., Style, Tone, and Audience), and incorporated detailed class definitions in the Response section. The third variation (SP3) refined the structured prompt by explicitly reiterating the expected labels in the Response section; this aims to further clarify the task, and enhance the consistency of the model’s outputs. All the prompts can be found in Appendix A.
2.3.3 Query Templates
Query templates (QT) define the structure in which citation sentences are presented to the model, both as examples during prompting and as queries during evaluation. Initially, we adopted a Simple Query template: the citation sentence was followed by “Class:”, either pre-filled with the correct class (in examples) or left blank (in queries). However, we observed that some models struggled to align with the expectation of responding solely with the class labels, leading to inconsistent performance and increased variance. To address this, we introduced a Multiple-choice Query template, where, following the citation sentence, the model is explicitly asked to identify the most likely citation intent, with the possible class labels presented as multiple-choice options. While this approach increases the token count per query, it demonstrates significant performance gains, as discussed in Section 3.2.
Dataset Rank Model Overall Zero-Shot One-Shot Few-Shot Many-Shot
SciCite 1 Qwen 2.5 – 14B 112 20 37 27 28
2 Mistral Nemo – 12B 22 0 0 12 10
3 Gemma 2 – 9B 4 0 3 1 0
4 Gemma 2 – 27B 2 0 0 0 2
ACL-ARC 1 Qwen 2.5 – 14B 24 4 4 8 8
2 Gemma 2 – 27B 4 0 4 0 0
Table 1: Model ranking based on Best-Performing Count (Overall & By Prompting Method).
Dataset Rank Model PM SP QT EM T F1-Score
SciCite 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.5 78.27
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 78.39
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 78.34
4 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.0 78.00
5 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.5 77.93
ACL-ARC 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 60.88
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.0 59.75
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 59.66
4 Qwen 2.5 – 14B One-shot SP3 Multiple-Choice Roles 0.5 64.87
5 Qwen 2.5 – 14B One-shot SP3 Multiple-Choice Roles 1.0 64.32
Table 2: Top-5 Configurations on SciCite and ACL-ARC (Q8).
2.3.4 Example Method Variations
In our experiments, we explored two methods for presenting examples to the models (EM): Inline and Roles.
In the Inline approach, example citation sentences and their corresponding classes (as per the defined query templates) are provided directly within the system prompt. After this setup, the prompt then transitions to the evaluation phase, where the “user” provides a citation sentence without a corresponding class and the model, acting as the “assistant”, predicts the correct class.
The Roles method, on the other hand, simulates a conversational exchange between the “user” and “assistant”. Each example follows a back-and-forth interaction, with the “user” providing a citation sentence and the “assistant” responding with the correct class. After several such examples, the interaction moves to the evaluation phase, where the “user” presents a new citation sentence, and the model predicts its class.
2.3.5 Temperature
Temperature (T) is a hyperparameter that controls the randomness or creativity of a language model’s outputs by adjusting the probability distribution of possible next tokens (Jurafsky and Martin, 2025). Lower temperatures (i.e., close to 00) correspond to greedy decoding, where the model deterministically selects the most probable token at each step, while higher temperatures (closer to $1$) introduce greater variability by allowing the model to sample from a broader range of options.
For classification tasks, we aim for the model to output the most probable class label. Therefore, we use a temperature of 00 as the baseline, ensuring fully deterministic predictions. To explore how controlled randomness affects the model’s behavior and classification performance, we also evaluate higher temperatures, such as 0.2, 0.5, and 1.0. A temperature of 0.2 introduces a small degree of randomness but still heavily favors the most probable answer, while 0.5 strikes a balance between randomness and determinism. At 1.0, randomness is maximized, allowing us to assess whether excessive stochasticity degrades performance.
2.4 Technical Specifications
We conducted our experiments on an M1 Max Mac Studio with $64$GB of memory, chosen to demonstrate the feasibility of running inference for Citation Intent Classification on commodity hardware.
For model hosting, we used LM Studio [^3] https://lmstudio.ai/ which offers an intuitive interface for testing and interacting with models hosted on HuggingFace or locally. It also supports a local server mode compatible with the OpenAI API, allowing interaction through an API accessible in multiple programming languages. A command-line interface (CLI) tool [^4] https://github.com/lmstudio-ai/lms is also available for managing the server without using the UI.
3 Experimental Results and Analysis
This section presents the results of our experiments, focusing on the performance of models across various configurations and parameter settings. We analyze the best-performing models, evaluate their rankings, and explore how specific parameters influence performance.
3.1 Model Performance over Configurations
Our first evaluation focuses on identifying the top-performing models across the configurations described in Section 2. We employed two complementary approaches: (i) the Best-Performer Evaluation identifies the single best model, and (ii) the Ranked Evaluation considers the relative performance of all models across configurations.
3.1.1 Best-Performer Evaluation
To identify the most effective models across configurations, we initially conducted a Best-Performer Evaluation. In particular, for each configuration, we examined a metrics table containing precision, recall, F1-score, and accuracy for all models. The model with the highest F1-score in each configuration was selected as the best-performing model – in case of a tie, we used Accuracy as the deciding factor.
The aggregated results of our evaluation are presented in Table 1; additionally, Table 2 lists the top-$5$ best-performing configurations on each dataset. It is evident that Qwen 2.5 14B was the most dominant model on both datasets, significantly outperforming others across all prompting methods.
Rank Model Score
1 Qwen 2.5 – 14B 124.164
2 Mistral Nemo – 12B 59.442
3 Gemma 2 – 27B 48.421
4 Gemma 2 – 9B 47.904
5 Qwen 2 – 7B 31.927
6 LLaMA 3.1 – 8B 23.591
7 LLaMA 3 – 8B 21.535
8 Phi 3.5 Mini – 3.8B 18.869
9 Phi 3 Medium – 14B 16.877
10 Gemma 2 - 2B 15.467
11 LLaMA 3.2 - 3B 14.603
12 LLaMA 3.2 - 1B 11.620
Table 3: Model ranking based on RRF on SciCite.
3.1.2 Ranked Evaluation
While the Best-Performer Evaluation provided insights into the most dominant models across configurations, it lacked the ability to capture nuances among high-performing models. For instance, a model that consistently ranked second in multiple configurations would not be reflected in that approach. To address this limitation, we adopted a ranking methodology inspired by Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). This approach allowed us to evaluate all $12$ models considering their relative performance across different configurations.
For each experimental configuration $c\in C$, models $M=\{m_{1},m_{2},\dots,m_{12}\}$ were ranked based on their F1-scores – in case of a tie, Accuracy was used to determine the rank. Each model was then assigned a score based on its rank, where the score $S(m_{k},c)$ for the $k$-th ranked model in configuration $c$ was defined as $S(m_{k},c)=\frac{1}{k}$. This score is inversely proportional to the rank, meaning that higher-ranked models (lower-$k$) receive higher scores, while lower-ranked models (higher-$k$) receive lower scores. To calculate the overall ranking score for each model $m$, we aggregated its scores across all configurations $c\in C$:
$Rank(m)=\sum_{c\in C}S(m_{k},c)$ (1)
Tables 3 and 4 present the results of this ranked evaluation. The ranked evaluation aligns with the Best-Performer Evaluation in identifying Qwen 2.5 14B as the most dominant model. However, it also offers insights into the relative performance of other models, highlighting distinctions among high performers.
3.2 Parameter Performance Analysis
In this section, we examine how different parameter configurations drive model performance. Identifying the most impactful parameters was not a straightforward task due to the complexity of the search space and the variety of configurations tested. To address this, we conducted a Quantile-Based Analysis, focusing on the top $5$% of configurations based on F1-scores. Examining the distribution of parameter settings within this high-performing subset, allows us to uncover trends and patterns that could guide the selection of optimal configurations.
For this analysis, we focused exclusively on the SciCite dataset, and the insights gained from it were then used to limit the range of experiments conducted on the ACL-ARC dataset. The findings from this analysis can be seen in Table 5 and are summarized below.
Rank Model Score
1 Qwen 2.5 – 14B 26.000
2 Gemma 2 – 27B 9.747
3 Gemma 2 – 9B 8.498
4 Qwen 2 – 7B 7.554
5 LLaMA 3.1 – 8B 6.437
6 Mistral Nemo – 12B 6.000
7 Phi 3.5 Mini – 3.8B 5.268
8 Phi 3 Medium - 14B 3.399
9 LLaMA 3.2 - 3B 3.281
10 LLaMA 3.2 - 1B 2.718
11 Gemma 2 - 2B 2.506
12 LLaMA 3 - 8B 2.396
Table 4: Model ranking based on RRF on ACL-ARC.
We discern that few-shot prompting was the most common prompting method among the top-performing configurations, accounting for nearly 45% of the results. Many-shot prompting followed closely with 34.83%, while one-shot and zero-shot prompting were significantly less represented, with only 16.85% and 3.37%, respectively. This suggests that providing several examples during prompting generally leads to better performance, with the optimal achieved at five examples (few-shot), followed by a slight decrease at larger numbers of examples.
In contrast, the temperature parameter exhibited a relatively even distribution among the top-performing configurations, with $0.0$ and $0.2$ being the most frequent settings (30.34% each). Higher temperatures, such as $1.0$ and $0.5$, were less common, suggesting that lower temperatures contribute to more consistent and deterministic outputs, which align better with the task requirements.
Similarly, the system prompt parameter was fairly evenly distributed among the top configurations, with SP3 being slightly more frequent (35.96%), followed by SP2 (32.58%), and SP1 (31.46%). While all system prompts contributed to high performance, these results suggest that SP3 has offered a slight advantage in effectiveness.
Parameter Setting Count Percent
Prompting Method Few-shot 40 44.94%
Many-shot 31 34.83%
One-shot 15 16.85%
Zero-shot 3 3.37%
Temperature 0.0 27 30.34%
0.2 27 30.34%
1.0 18 20.22%
0.5 17 19.10%
System Prompt SP3 32 35.96%
SP2 29 32.58%
SP1 28 31.46%
Query Template Multiple-Choice 83 93.26%
Simple 6 6.74%
Examples Method Roles 44 51.16%
Inline 42 48.84%
Table 5: Parameter Performance Analysis of Top 5% Configurations on SciCite (sorted by percentage).
The query template parameter showed a strong preference for multiple-choice templates, which appeared in 93.26% of the top-performing configurations. This overwhelming dominance suggests that multiple-choice queries are significantly more effective for the evaluated task, likely due to their structured.
The examples method parameter showed a near-equal split between roles (51.16%) and inline (48.84%), indicating that both approaches are effective, with no clear advantage for either method among the top-performing configurations.
These findings offer valuable insights into the parameter settings that consistently enhance performance. To verify this analysis, we also performed a Chi-Square statistical test, outlined in Appendix B.
4 Fine-tuning
In this section, we investigate the impact of fine-tuning in the the performance of the instruction-tuned Qwen 2.5 14B model on the task of citation intent classification.
Parameter Value
Learning Rate 5e-5
Epochs 10
Batch Size 16
Cutoff Length 512
Optimizer AdamW
Compute Type fp16
Warmup Steps 500
DeepSpeed Offload Enabled
DeepSpeed Stage 3
LoRA Rank 8
LoRA Alpha 16
LoRA Dropout 0.1
Table 6: Training Parameters for Fine-Tuning Qwen 2.5 14B.
4.1 Training Configuration
For this experiment, we used the SciCite and ACL-ARC datasets, which were converted into the Alpaca format (Taori et al., 2023). This format includes a system prompt, an instruction, the citing sentence, and the true label. The use of Supervised Fine-Tuning (SFT) was motivated by its ability to adapt pre-trained models to specific tasks using minimal labeled data, making it an effective approach for citation intent classification.
To fine-tune the Qwen 2.5 14B Instruct model, we used LLaMA-Factory (Zheng et al., 2024) on an AWS g6e.12xlarge EC2 instance equipped with 4 NVIDIA L40S GPUs, providing a combined GPU memory of 192GB. The training process was conducted using fp16 mixed precision to optimize memory usage and speed. To prevent memory issues during fine-tuning, we employed DeepSpeed ZeRO Stage 3 Offload (Ren et al., 2021), which enabled efficient memory management by offloading optimizer states and gradients to the CPU. The training parameters included a learning rate of 5e-5, a batch size of 16, and 10 epochs. The model was optimized using AdamW, with a cutoff length of 512 tokens for input sequences.
To optimize the fine-tuning process, we employed Low-Rank Adaptation (LoRA), a widely used parameter-efficient fine-tuning (PEFT) method (Hu et al., 2021; Han et al., 2024). LoRA enables efficient adaptation of large language models by freezing the pre-trained model weights and introducing trainable low-rank matrices into the Transformer layers, significantly reducing the number of trainable parameters while maintaining performance. For this experiment, we configured LoRA with a rank of 8, alpha of 16, and a dropout of 0.1, which allowed us to fine-tune the model effectively without exceeding memory constraints. The outlined training parameters are also summarized in Table 6.
4.2 Results
After fine-tuning, we evaluated the new model using the same prompting methods as in Section 2.3.1 (i.e., zero-shot, one-shot, few-shot, and many-shot) to draw comparisons with the baseline Qwen 2.5 14B Instruct model. Although in the previous evaluation we only used 8-bit quantized models (Q8) due to the extensiveness of our experiment space, in this experiment, we also evaluate the 16-bit floating-point (FP16) version of the model to examine the impact of bit precision on performance.
Model SciCite ACL-ARC
Instruct Q8 78.33 61.04
Instruct FP16 80.41 61.73
Fine-tuned Q8 86.47 68.48
Fine-tuned FP16 86.84 67.73
Table 7: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants on the SciCite and ACL-ARC datasets.
The results in Table 7 show that fine-tuning improved performance on both SciCite and ACL-ARC. On SciCite, the fine-tuned FP16 model achieved an F1-score of 86.84, an 8% improvement over the instruction-tuned baseline, while the fine-tuned Q8 model achieved a comparable score of 86.47. On ACL-ARC, the fine-tuned Q8 model achieved the highest F1-score of 68.48, outperforming the instruction-tuned baseline by nearly 11%.
Precision settings showed minor differences, with FP16 models generally outperforming Q8 on SciCite, while Q8 slightly outperformed FP16 on ACL-ARC. This suggests that while FP16 precision is generally advantageous, Q8 remains competitive and may offer robustness for tasks with more granular classification schemes, such as ACL-ARC’s 6-class setup.
Table 8 highlights the impact of fine-tuning across prompting methods. On SciCite, fine-tuning led to consistent improvements, particularly in zero-shot and one-shot settings, where the fine-tuned Q8 model improved by nearly 10% over the instruction-tuned baseline. In few-shot and many-shot setups, performance approached saturation, with the FP16 model achieving the highest F1-scores of 86.84 and 85.79, respectively. On ACL-ARC, fine-tuned models showed similar trends, with the Q8 model achieving the highest F1-scores in one-shot (68.48) and few-shot (64.35) setups, outperforming FP16 in most cases. These results demonstrate that fine-tuning enhances generalization in low-context scenarios while effectively leveraging additional context during inference.
Dataset Model Zero- One- Few- Many-
Shot Shot Shot Shot
SciCite Instruct Q8 75.26 76.29 78.33 78.32
Instruct FP16 75.38 77.22 80.41 78.94
Fine-tuned Q8 84.84 85.46 86.47 85.62
Fine-tuned FP16 84.49 85.38 86.84 85.79
ACL-ARC Instruct Q8 50.02 60.02 61.04 59.75
Instruct FP16 50.17 61.73 59.04 54.46
Fine-tuned Q8 59.40 68.48 64.35 65.43
Fine-tuned FP16 60.00 67.73 64.34 64.62
Table 8: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants on the SciCite and ACL-ARC datasets divided by prompting method.
4.3 Discussion
The primary goal of this work was to investigate the capability of large language models to perform citation intent classification, rather than to compete directly with state-of-the-art methods. Nonetheless, it is worth noting that the fine-tuned model achieved performance on SciCite within less than a 3% margin of the best-reported results in the literature, surpassing most of the existing approaches (see Table 9 - Section 5). This demonstrates that LLMs can approach the performance of specialized models, even without task-specific architectures or optimization.
We consider a key advantage of LLMs to be their ease of use and deployment. Tools such as LM Studio and Ollama [^5] https://ollama.com/ allow models like those outlined in this paper to be deployed locally with zero technical expertise, making them accessible to users without a computer science background. In addition to their accessibility, LLMs offer significant adaptability. Unlike traditional methods, which often require complex pretraining or domain-specific tuning, LLMs can be fine-tuned for a wide range of scientometric tasks, such as citation recommendation, paper summarization, or trend analysis, without the need for bespoke architectures. Furthermore, LLMs can scale effectively to new scientific domains or datasets, requiring only small amounts of task-specific data to adapt to underexplored scientific fields. For instance, our fine-tuned models were trained only on the several thousand citing sentences provided by our task-specific datasets, whereas SciBERT has been trained on millions of scientific papers.
The promising results observed in this study suggest that the performance of LLMs in citation intent classification can be further improved with techniques such as chain-of-thought prompting and reasoning-focused models. These methods could enhance the models’ ability to better distinguish subtle differences in citation intent, refining their predictions and improving overall classification accuracy.
5 Related Work
The theoretical beginnings of citation analysis can be traced back to foundational works such as Garfield et al.’s (1965) identification of reasons for citation Moravcsik and Murugesan’s (1975) studies on citation function. Early annotation schemes, such as those by Spiegel-Rosing (1977), were later adapted by Teufel et al. (2006) for supervised machine learning approaches to citation classification.
Jurgens et al. (2018) introduced the ACL-ARC dataset, which contains nearly $2,000$ citations from papers in the NLP field, annotated for their function with a classification scheme of six classes. Pride and Knoth (2020) extended this classification scheme by refining the comparison class to capture similarities, differences, and disagreement. Around the same time, Cohan et al. (2019) proposed a multitask model incorporating structural information from scientific papers. They also introduced the SciCite dataset, which is significantly larger and spans multiple scientific domains, with three intent classes.
Beltagy et al. (2019) introduced SciBERT, a BERT-based (Devlin et al., 2019) pre-trained language model (PLM) for scientific text, which has since become the backbone of many citation intent classification methods. SciBERT has been widely adopted due to its ability to generalize across scientific domains. For example, Mercier et al. (2021) introduced ImpactCite, an XLNet-based method for citation impact analysis, which was later used by Paolini et al. (2024) to achieve state-of-the-art results on the SciCite dataset. Paolini et al. demonstrated the effectiveness of ensemble classifiers combining fine-tuned SciBERT and XLNet models. On the ACL-ARC dataset, Hu et al. (2022) achieved state-of-the-art results by studying domain adaptation of language models using a variational masked autoencoder.
Method SciCite ACL-ARC
Feature-rich RF (Jurgens et al., 2018) 53.00
Structural Scaffolds (Cohan et al., 2019) 84.00 67.90
SciBERT (Beltagy et al., 2019) 85.22
ImpactCite (Mercier et al., 2021) 88.93
VarMAE Hu et al. (2022) 86.32 76.50
CitePrompt (Lahiri et al., 2023) 86.33 68.39
EnsIntWS (Paolini et al., 2024) 89.46
EnsIntWoS (Paolini et al., 2024) 88.48
MTL Finetuning (Search) (Shui et al., 2024) 85.25 64.56
MTL Finetuning (TRL) (Shui et al., 2024) 85.35 75.57
Table 9: Reported F1-Scores of notable works from the literature, sorted chronologically.
Recent research has continued to explore PLM-based methods for citation intent classification. Lahiri et al. (2023) used a prompt-based learning approach on SciBERT to identify citation intent, while Shui et al. (2024) propose a multi-task learning framework that jointly fine-tunes SciBERT on a dataset of primary interest together with multiple auxiliary datasets to take advantage of additional supervision signals. Kunnath et al. (2023) explored various prompting and tuning strategies on SciBERT, including fixed and dynamic context prompts, and found that parameter updating with prompts improved performance. They also briefly experiment on LLMs by evaluating the zero-shot performance of GPT-3.5, which performed well on their recently introduced ACT2 dataset but poorly on the ACL-ARC dataset. However, GPT-3.5 was not evaluated on SciCite. In contrast, our work is the first to focus entirely on evaluating and fine-tuning numerous open-weight LLMs without using models pre-trained specifically for the scientific domain.
Table 9 provides a summary of the reported F1-scores of notable works in the literature, highlighting the progression of methods and datasets. SciBERT-based methods dominate the field, while our work is the first to examine LLMs on this task without any reliance to SciBERT.
6 Conclusions
This study explored the capabilities of open LLMs for citation intent classification, focusing on their performance with and without task-specific training. We found that while instruction-tuned LLMs perform reasonably well through in-context learning, supervised fine-tuning with minimal data leads to significant improvements. Additionally, our experiments on prompting-related parameters offer practical insights into optimizing model performance. To support further research, we provide our evaluation platform and models to the research community.
7 Limitations
While our study provides valuable insights into the performance of LLMs on the task of Citation Intent Classification, it is not without limitations. First, it was not feasible to evaluate all recent models comprehensively due to the rapid pace of advancements in the field. To address this, we plan to update our GitHub repository with any notable new models as they emerge. Second, the study did not consider the time, energy consumption, and compute costs associated with running LLMs compared to traditional methods, as this was deemed out of scope; instead, we focused solely on assessing the viability of the approach. Finally, due to computational constraints, our analysis was limited to only the small and medium parameter counts of the selected models (1-27B parameters).
Acknowledgments
This work was co-funded by the EU Horizon Europe projects SciLake (GA: 101058573) and GraspOS (GA: 101095129). Part of this work utilized Amazon’s cloud computing services, which were made available via GRNET under the OCRE Cloud framework, providing Amazon Web Services for the Greek Academic and Research Community. The authors wish to also thank Dimitris Roussis and Leon Voukoutis, members of Athena RC’s Institute for Speech and Language Processing for their assistance on key points of this work.
Appendix A Appendix – System Prompts, Query Templates, Example Methods
In this appendix, we present the System Prompts used in our experiments as defined in Section 2.3.2. We also supply examples of the Query Templates and Example Method Variations outlined in Sections 2.3.3 and 2.3.4. Listings 12, and 3 show the system prompts used for SciCite. The prompts for ACL-ARC follow the same structure, with the appropriate changes to the class definitions and can be found in the supplementary materials. Listings 4 and 5 show examples of the query templates we defined and Listings 6 and 7 show the different ways that we supplied the samples in our prompts.
You are an expert researcher tasked with classifying the intent of a citation in a scientific publication.
The three (3) possible classes are the following: "background information", "method", "results comparison".
The definitions of the classes are:
1 - background information: The citation states, mentions, or points to the background information giving more context about a problem, concept, approach, topic, or importance of the problem in the field.
2 - method: Making use of a method, tool, approach or dataset.
3 - results comparison: Comparison of the papers results/findings with the results/findings of other work.
For each given sentence, you must analyse only the citation with the @@CITATION@@. You must assign only one class to each citation. Only return the class name, with no elaboration.
Listing 1: SP1 for SciCite
# CONTEXT #
You are an expert researcher tasked with classifying the intent of a citation in a scientific publication.
# OBJECTIVE #
You will be given a sentence containing a citation, you must output the appropriate class as an answer.
# CLASS DEFINITIONS #
The three (3) possible classes are the following: "background information", "method", "results comparison".
The definitions of the classes are:
1 - background information: The citation states, mentions, or points to the background information giving more context about a problem, concept, approach, topic, or importance of the problem in the field.
2 - method: Making use of a method, tool, approach or dataset.
3 - results comparison: Comparison of the papers results/findings with the results/findings of other work.
# RESPONSE RULES #
You must strictly adhere to the following rules for your response:
- For each sentence, you must analyse only the citation with the @@CITATION@@ tag.
- You must assign only one class to each citation.
- Only respond with the class name, with no explanation or elaboration.
- Only answer with one or two words.
- Always be very brief.
Listing 2: SP2 for SciCite
# CONTEXT #
You are an expert researcher tasked with classifying the intent of a citation in a scientific publication.
# OBJECTIVE #
You will be given a sentence containing a citation. You must classify the intent of the citation by assigning it to one of three predefined classes.
# CLASS DEFINITIONS #
The three (3) possible classes are the following: "background information", "method", "results comparison".
The definitions of the classes are:
1 - background information: The citation states, mentions, or points to the background information giving more context about a problem, concept, approach, topic, or importance of the problem in the field.
2 - method: Making use of a method, tool, approach or dataset.
3 - results comparison: Comparison of the papers results/findings with the results/findings of other work.
# RESPONSE RULES #
- Analyze only the citation marked with the @@CITATION@@ tag.
- Assign exactly one class to each citation.
- Respond only with the exact name of one of the following classes: "background information", "method", or "results comparison".
- Do not provide any explanation or elaboration.
Listing 3: SP3 for SciCite
These results are in contrast with the findings of Santos et al.@@CITATION@@, who reported a significant association between low sedentary time and healthy CVF among Portuguese.
Class: results comparison
Listing 4: Example of a Simple Query Template (QT1)
These results are in contrast with the findings of Santos et al.@@CITATION@@, who reported a significant association between low sedentary time and healthy CVF among Portuguese.
### Question: Which is the most likely intent for this citation?
a) background information
b) method
c) results comparison
### Answer: results comparison
Listing 5: Example of a Multiple-Choice Query Template (QT2)
[Start of system prompt]
[system prompt content]
# EXAMPLES #
These results are in contrast with the findings of Santos et al.@@CITATION@@, who reported a significant association between low sedentary time and healthy CVF among Portuguese.
Class: results comparison
Two or fewer cancer-containing cores has been shown to correlate with tumour volume in US series (@@CITATION@@).
Class: background information
[end of system prompt]
role: user
message: The regions of dhfr and dhps genes containing the mutations for antifolate resistance were amplified as described elsewhere (@@CITATION@@) and sequenced for detection of the mutations..
role: assistant
message: [prediction]
Listing 6: Example of the “Inline” Examples Method (EM1)
[Start of system prompt]
[system prompt content]
[end of system prompt]
role: user
message: These results are in contrast with the findings of Santos et al.@@CITATION@@, who reported a significant association between low sedentary time and healthy CVF among Portuguese.
role: assistant
message: results comparison
role: user
message: Two or fewer cancer-containing cores has been shown to correlate with tumour volume in US series (@@CITATION@@).
role: assistant
message: background information
role: user
message: The regions of dhfr and dhps genes containing the mutations for antifolate resistance were amplified as described elsewhere (@@CITATION@@) and sequenced for detection of the mutations..
role: assistant
message: [prediction]
Listing 7: Example of the “Roles” Examples Method (EM2)
Appendix B Chi-Square Statistical Test
To complement the parameter performance analysis outlined in Section 3.2, we also conducted a Chi-Square Test of Independence to evaluate the relationship between parameter settings and F1-scores. This approach allowed us to determine whether specific parameter settings were significantly associated with performance outcomes and validate the results of our primary analysis. The results of the test are summarized in Table 10.
The test revealed several significant relationships between parameter settings and performance outcomes. Specifically, the Prompting Method ($\chi^{2}=22.46$, $p=5.23e-05$), System Prompt ($\chi^{2}=14.97$, $p=0.00056$), and Query Template ($\chi^{2}=41.86$, $p=9.79e-11$) were all found to have statistically significant associations with performance, indicating that these parameters play a critical role in influencing F1-scores. In contrast, Temperature ($\chi^{2}=4.31$, $p=0.230$) and Examples Method ($\chi^{2}=0.01$, $p=0.911$) did not show significant relationships, suggesting that these parameters may not independently affect performance outcomes. These results are aligned with the findings of our Quantile-based analysis.
Parameter Chi-Square ($\chi^{2}$) p-value Significance Level
Prompting Method 22.46 5.23e-05 ***
Temperature 4.31 0.230
System Prompt 14.97 0.00056 ***
Query Template 41.86 9.79e-11 ***
Examples Method 0.01 0.911
Table 10: Chi-Square Test of Independence results for parameter settings and F1-Scores. (Significance levels: ***p < 0.001, **p < 0.01, p < 0.05)
Appendix C Appendix – Models
Table 11 outlines the models used in our evaluation (Section 2.1) along with their context lengths and HuggingFace links.
Model Context Length HF Links
Llama 3.2 Instruct – 1B 128K hugging-quants/Llama-3.2-1B-Instruct-Q8_0-GGUF
Llama 3.2 Instruct – 3B 128K hugging-quants/Llama-3.2-3B-Instruct-Q8_0-GGUF
Llama 3.1 Instruct – 8B 128K lmstudio-community/Meta-Llama-3.1-8B-Instruct-GGUF
Llama 3 Instruct – 8B 8192 lmstudio-community/Meta-Llama-3-8B-Instruct-GGUF
Mistral Nemo Instruct – 12B 128K lmstudio-community/Mistral-Nemo-Instruct-2407-GGUF
Phi3 Medium Instruct – 14B 128K bartowski/Phi-3-medium-128k-instruct-GGUF
Phi3.5 Mini Instruct – 3.8B 128K bartowski/Phi-3.5-mini-instruct-GGUF
Gemma 2 Instruct – 2B 8192 lmstudio-community/gemma-2-2b-it-GGUF
Gemma 2 Instruct – 9B 8192 bartowski/gemma-2-9b-it-GGUF
Gemma 2 Instruct – 27B 8192 bartowski/gemma-2-27b-it-GGUF
Qwen 2 Instruct – 7B 32K Qwen/Qwen2-7B-Instruct
Qwen 2.5 Instruct – 14B 32K lmstudio-community/Qwen2.5-14B-Instruct-GGUF
Table 11: The models used in our evaluation.
Appendix D Appendix – Top Performing Configurations
Table 12 outlines the F1-scores of the best performing configuration for each model. Table 13 shows the Top 20 configurations of our evaluation for both datasets.
Dataset Model PM SP QT EM T F1-Score
SciCite Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.5 78.33
Mistral Nemo – 12B Few-shot SP2 Multiple-Choice Inline 0.0 77.38
Gemma 2 – 27B Few-shot SP3 Multiple-Choice Roles 0.2 76.16
Gemma 2 – 9B Few-shot SP3 Multiple-Choice Roles 0.2 74.97
Phi 3 Medium – 14B Many-shot SP3 Multiple-Choice Roles 0.2 74.67
LLaMA 3 – 8B Many-shot SP1 Multiple-Choice Roles 0.0 74.39
Qwen 2 – 7B Few-shot SP2 Simple Roles 0.2 72.89
LLaMA 3.1 – 8B Few-shot SP1 Simple Roles 0.0 72.46
Gemma 2 - 2B Many-shot SP3 Multiple-Choice Roles 0.5 68.79
Phi 3.5 Mini – 3.8B One-shot SP3 Multiple-Choice Roles 0.2 68.25
LLaMA 3.2 - 3B Many-shot SP3 Multiple-Choice Roles 0.2 67.99
LLaMA 3.2 - 1B Few-shot SP1 Simple Roles 0.2 45.44
ACL-ARC Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 61.04
Gemma 2 – 9B Few-shot SP3 Multiple-Choice Roles 1.0 57.19
Gemma 2 – 27B One-shot SP3 Multiple-Choice Roles 1.0 57.00
Mistral Nemo – 12B One-shot SP3 Multiple-Choice Inline 0.2 48.11
Qwen 2 – 7B Few-shot SP3 Multiple-Choice Roles 0.0 47.50
LLaMA 3.1 – 8B Many-shot SP3 Multiple-Choice Roles 0.2 46.43
Phi 3.5 Mini – 3.8B Zero-shot SP3 Multiple-Choice 0.5 41.82
Phi 3 Medium - 14B One-shot SP3 Multiple-Choice Roles 0.2 39.30
LLaMA 3.2 - 3B One-shot SP2 Multiple-Choice Roles 0.2 37.82
LLaMA 3 - 8B Zero-shot SP2 Multiple-Choice 0.2 37.29
Gemma 2 - 2B Few-shot SP2 Multiple-Choice Roles 0.0 36.80
LLaMA 3.2 - 1B Many-shot SP2 Multiple-Choice Roles 0.5 24.60
Table 12: F1-Scores of the top-performing configuration for each model.
Dataset Rank Model PM SP QT EM T Precision Recall F1-Score Accuracy
SciCite 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.5 78.27 78.91 78.33 80.71
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 78.39 78.65 78.31 80.71
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 78.34 78.65 78.25 80.66
4 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.0 78.00 79.40 78.23 80.76
5 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.5 77.93 79.29 78.13 80.66
6 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.2 77.90 79.22 78.09 80.66
7 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.0 78.07 78.57 78.07 80.49
8 Qwen 2.5 – 14B Few-shot SP2 Multiple-Choice Roles 0.0 77.63 78.95 78.02 80.28
9 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 1.0 77.67 79.29 77.98 80.33
10 Qwen 2.5 – 14B Few-shot SP2 Multiple-Choice Roles 1.0 77.60 78.51 77.86 80.23
11 Qwen 2.5 – 14B Few-shot SP2 Multiple-Choice Roles 0.2 77.35 78.73 77.77 80.06
12 Qwen 2.5 – 14B Many-shot SP2 Multiple-Choice Roles 1.0 77.15 79.05 77.72 80.23
13 Qwen 2.5 – 14B Few-shot SP2 Multiple-Choice Roles 0.5 77.37 78.55 77.67 79.96
14 Qwen 2.5 – 14B Many-shot SP2 Multiple-Choice Roles 0.2 77.14 78.97 77.63 80.12
15 Qwen 2.5 – 14B Many-shot SP2 Multiple-Choice Roles 0.0 77.14 78.80 77.55 80.06
16 Qwen 2.5 – 14B Many-shot SP2 Multiple-Choice Roles 0.5 77.04 78.84 77.52 80.06
18 Mistral Nemo – 12B Few-shot SP1 Simple Inline 0.2 79.11 76.05 77.39 79.10
17 Mistral Nemo – 12B Few-shot SP2 Multiple-Choice Inline 0.0 78.20 76.78 77.38 79.58
19 Mistral Nemo – 12B Few-shot SP1 Simple Inline 0.0 79.29 75.90 77.33 79.26
20 Mistral Nemo – 12B Few-shot SP1 Multiple-Choice Inline 0.0 76.80 78.55 77.30 78.45
ACL-ARC 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 60.88 65.36 61.04 66.91
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.0 59.75 64.49 60.24 65.47
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 59.66 64.49 60.07 65.47
4 Qwen 2.5 – 14B one-shot SP3 Multiple-Choice Roles 0.5 64.87 57.80 60.02 66.19
5 Qwen 2.5 – 14B one-shot SP3 Multiple-Choice Roles 1.0 64.32 58.44 59.83 66.91
6 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Inline 0.0 63.34 62.35 59.75 66.91
7 Qwen 2.5 – 14B one-shot SP3 Multiple-Choice Roles 0.0 64.49 57.16 59.55 65.47
8 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Inline 0.5 62.97 62.12 59.40 66.19
9 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Inline 0.2 62.98 61.71 59.30 66.19
10 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Inline 0.5 63.32 62.29 59.24 65.47
11 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Inline 0.0 62.96 61.65 58.82 64.75
12 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Inline 1.0 60.10 60.93 58.68 65.47
13 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.5 58.54 61.80 58.50 65.47
14 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Inline 1.0 60.78 61.65 58.39 64.75
15 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Inline 0.2 62.27 60.98 58.01 64.03
16 Gemma 2 – 9B Few-shot SP3 Multiple-Choice Roles 1.0 57.59 58.33 57.19 62.59
17 Gemma 2 – 27B one-shot SP3 Multiple-Choice Roles 1.0 57.16 59.59 57.00 61.15
18 Gemma 2 – 27B one-shot SP3 Multiple-Choice Inline 0.0 57.50 58.11 56.55 60.43
19 Gemma 2 – 27B one-shot SP3 Multiple-Choice Inline 0.2 57.50 58.11 56.55 60.43
20 Qwen 2.5 – 14B one-shot SP3 Multiple-Choice Roles 0.2 61.92 54.06 56.54 65.47
Table 13: Top-20 Configurations on SciCite and ACL-ARC (Q8).